The measures of two complementary angles are in the ratio 1:5.Find the measure of each angle.
Found 2 solutions by Theo, amoresroy:
Answer by Theo(13342)

if the angles are complementary, then their sum is 90 degrees.
let x = one of the angles and let y = the other.
you get:
x + y = 90
one of the angles is in the ratio of 1:5 with the other angle.
let x be that angle.
ratio formula says x/y = 1/5
solve for x to get x = (1/5)*y
replace y in the formula of x + y = 90 to get:
(1/5)*y + y = 90
combine like terms to get (6/5)*y = 90
multiply both sides of equation by (5/6) to get:
y = (5/6) * 90 = 75 degrees.
since x + y = 90, then x must be 15 degrees.
x = 15
y = 75
x/y = 15/75 = 1/5
looks good.
the angles measure 15 degrees and 75 degrees.


Answer by amoresroy(361)

Let x = the measure of smaller angle
5x = the measure of bigger angle
Complementary angles add up to 90
x+5x = 90
6x = 90
x = 15
5x = 5(15) = 75